OpenStudy (anonymous):

Solve by elimination 0.3y-0.2y=4 0.4x+0.5y=-29/19 answer in fraction form

OpenStudy (saifoo.khan):

fractin version?

OpenStudy (anonymous):

answer needs to be in a fraction

OpenStudy (saifoo.khan):

-2335/209, 1230/209

OpenStudy (anonymous):

you do these fast!

OpenStudy (saifoo.khan):

:D

OpenStudy (anonymous):

ok that is incorrect what is the interger version?

OpenStudy (saifoo.khan):

decimal

OpenStudy (anonymous):

yes please

OpenStudy (saifoo.khan):

hey u typd the wrong ques~ :O chk the 1st eq!

OpenStudy (anonymous):

oh yes 0.3x-0.2y=4 0.4x+0.5y=-29/19

OpenStudy (saifoo.khan):

140/19, -170/19

OpenStudy (anonymous):

thank you